OpenStudy (anonymous):

I really need help! Samantha and Jake are training for a marathon. To prepare for this marathon they have been training and tracking their progress periodically. In the first week of training Jake ran an average of 7.75 minutes per mile. Later, in week four of training he ran an average of 7.00 minutes per mile. In the first week of training Samantha ran an average of 8.10 minutes per mile. Later, in week six of training she ran an average of 6.60 minutes per mile. Assuming that Jake and Samantha continue to train and improve their times at the same rate your task is to determine which

OpenStudy (anonymous):

which week they will have the same average minutes per mile. To complete this task follow the steps below. 1. Determine the equation of a line in standard form that represents Jake’s training progress. His progress corresponds to the points (1, 7.75) and (4, 7.00). 2. Determine the equation of a line in standard form that represents Samantha’s training progress. Her progress corresponds to the points (1, 8.10) and (6, 6.60) 3. Solve the system of equations (you must show all your work to receive full credit). 4. In which week will Jake and Samantha have the same average minutes per mile?

OpenStudy (anonymous):

Step 1: Y-7.75 = - 0.75/3 (X - 1) Step 2: Y-8.10 = - 1.50/5 (X - 1)

OpenStudy (anonymous):

1. Jake progress: x + 4y = 32 2. Samantha progress: 3x + 10y = 84 3. Solve ( 8, 6) 4. Week 6, at 6 min/ mile

OpenStudy (anonymous):

Woww thank you so much! :) I really, really appreciate it!

OpenStudy (anonymous):

How did you solve the system of equations? Would you mind showing me?

OpenStudy (anonymous):

1. Jake progress: m = ( 7.75 - 7 ) / ( 1 -4) = -.25 -> y - 7 = -.25 ( x -4)

OpenStudy (anonymous):

Use elimination method

OpenStudy (anonymous):

I don't know that method D: I kind of suck at algII sooo